OpenStudy (anonymous):

The sum of Pete's and Sam's ages is 30. Five years ago, Pete was 3 times as old as Sam. How old is Sam? Let P = Pete's age, S = Sam's age, and P + S = 30. Which of the following equations would complete the system?

OpenStudy (anonymous):

P + S = 30 Five years ago, Peter's age : (P - 5) Sam's Age is: (S - 5) Now, (P - 5) = 3(S - 5) P - 5 = 3S - 15 P - 3S = -10

OpenStudy (anonymous):

Which of the following equations would complete the system? Where are the following equations?

OpenStudy (radar):

S+P=30 P-5=3(S-5) P-5=3S-15 P-3S=-10 Subtract P+S=30 P-3S=-10 -P-S =-30 ----------- -4s=-40 S=10 P=20

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

You first give the choices..

OpenStudy (anonymous):

A) P-5=3S B) P-5=3S-5 C) P-5=3S-15

OpenStudy (anonymous):

P - 5 = 3S - 15 See my solution did you get this equation there??

OpenStudy (anonymous):

Check it in the second last line..

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

John is twice as old as Mary. The sum of their ages is 21. How old is Mary? Let J = John's age and M = Mary's age.

OpenStudy (anonymous):

J = 2M J + M = 21 2M + M = 21 3M = 21 Find M from here.

OpenStudy (zzr0ck3r):

today is not my day, sorry:)

OpenStudy (anonymous):

Ha ha ha ha.. Just concentrate..

OpenStudy (anonymous):

What you get for M??

OpenStudy (anonymous):

i dont even know how to solve this problem.

OpenStudy (anonymous):

See John is twice as old as Mary. Mary is M years old, then John will how many years old??

OpenStudy (anonymous):

im looking for the equation that represents the problem.

OpenStudy (anonymous):

J = 2M and the other equation is: J + M = 21

OpenStudy (anonymous):

The sum of two numbers is 44, and the larger number is 2 more than the smaller number. What is the smaller number? If x = the smaller number and y = the larger number, then which of the following systems of equations represents the word problem?